# Information flow-enhanced precision in collisional quantum thermometry

### v2

@@ IV Assessing information flow and its role in the thermometric task
Conclusion.– We have built a model to improve and increase the sensitivity of a quantum thermometer based on the use of a multilayered system. Even with a short interaction time, i.e. without requesting thermalization of the system with the thermometer, we can adjust the sensitivity of the latter by increasing the number of its ancillary chains. Our model has proven to be effective when we use only one ancilla in each chain, i.e. j=1, as long as we put a sufficient number of chains $k$. However, although this means fewer iterations, any increase in the number of chains $k$ will result in an exponential increase in the Hilbert space of the system, computationally limiting the construction of the sensor. We have unveiled a tight link between the sensitivity of the estimation of temperature as provided by the QFI and the information flow among various parts of the compound. The accumulation of information entailed by the information flow leads to better sensitivity (i.e. a larger QFI). This can also be seen in the increase of mutual information with the number of chains comprised in the thermometer.

### v3

@@ IV Assessing information flow and its role in the thermometric task
V Conclusion
We have built a model to improve and increase the sensitivity of a quantum thermometer based on the use of a multilayered system. Even with a short interaction time, i.e. without requesting thermalization of the system with the thermometer, we can adjust the sensitivity of the latter by increasing the number of its ancillary chains. Our model has been shown to be effective when we use only one ancilla in each chain, i.e. $j=1$, as long as we put a sufficient number of chains $k$. However, although this means fewer iterations, any increase in the number of chains $k$ will result in an exponential increase in the Hilbert space of the system, computationally limiting the construction of the sensor. We have unveiled a tight link between the sensitivity of the estimation of temperature as provided by the QFI and the information flow among various parts of the compound. The accumulation of information entailed by the information flow leads to better sensitivity (i.e. a larger QFI). This can also be seen in the increase of mutual information with the number of chains comprised in the thermometer. Note, however, that the relationship between information flow and the improvement of a sensor’s sensitivity also depends on the type of interaction between the system and the probe. To understand whether this connection is valid, it would be necessary to perform tests on various sensor models with different interaction dynamics. Another point that can also be worked on in the improvement of sensors is the use of other quantum resources such as coherence and entanglement. In our sensor, we measured both, but we did not find the creation of coherence in the dynamics of the ancillae qubits. As for entanglement, we measured the quantum negativity, which, although its maximum increases in the same way as the maximum of the QFI, has a small influence on the dynamics, having a value of 0.065QFImax.